OpenStudy (anonymous):

How many kilojoules of energy are absorbed when 800.0g of water is heated from 20.0º C to 50.0º C?

OpenStudy (chmvijay):

specific heat of water do you know they have not mentioned !!!

OpenStudy (anonymous):

specific heat of water is 40184j/g celsious

OpenStudy (anonymous):

4.184

OpenStudy (chmvijay):

Q=M*S*Delta T put all the values in this equation m for mass S for specific heat of water and Delta T is difference in temperature

OpenStudy (anonymous):

100416

OpenStudy (chmvijay):

what is that

OpenStudy (anonymous):

that is the answer i got when i

OpenStudy (anonymous):

(800.0)(4.184)(30.0)

OpenStudy (anonymous):

How many kilojoules of energy are absorbed when 800.0g of water is heated from 20.0º C to 50.0º C? Question 3 options: 1) 39.0 kJ 2) 163 kJ 3) 100 kJ 4) 62.8 kJ

OpenStudy (anonymous):

how am i adding in kilojoules/?

OpenStudy (anonymous):

don't i need to convert to kilojoules??

OpenStudy (chmvijay):

yaaa the unit for 4.184 is joules the answer you got is in joules that is 100416 J 1000J=1Kj 100416J=100416/1000=100.416KJ

OpenStudy (chmvijay):

the answer is 3rd one

OpenStudy (anonymous):

ok..just so i understand. To convert into kj, just divide my final answer by 1000 ( which is the 1kj=1000):?

OpenStudy (chmvijay):

yaaa dude

OpenStudy (anonymous):

heat difference is the delta t right?

OpenStudy (chmvijay):

yaaaa delta T =T2-T1

OpenStudy (chmvijay):

its temperature difference not heat difference

OpenStudy (anonymous):

ok...mind if you watch while i walk through another question?

OpenStudy (anonymous):

i learn best by return demonstration

OpenStudy (anonymous):

400.0g of a metal absorbs 10000. J of heat energy and its temperature rises from 20.0º C to 103.0º C. What is the specific heat of the metal? Question 4 options: 1) 0.301 J/gº C 2) 0.255 J/gº C 3) 3.32 J/g º C 4) 0.243 J/g ºC

OpenStudy (chmvijay):

find out S all other have given in this

OpenStudy (anonymous):

so i'm using the specific heat formula...heat/massxdelta t

OpenStudy (chmvijay):

yaaa right

OpenStudy (anonymous):

mass..400.0 delta t =20-103=83...which is my heat?

OpenStudy (anonymous):

1000?

OpenStudy (chmvijay):

what 1000? mean

OpenStudy (anonymous):

is that my heat?

OpenStudy (chmvijay):

S=q/Delta T*M =10000/83*400 =0.3J/g C

OpenStudy (chmvijay):

no its 10000J right why 1000J u want to take

OpenStudy (anonymous):

oh yea..sorry...didn't add the extra zero..

OpenStudy (chmvijay):

LOL!!!!

OpenStudy (anonymous):

are you going to be online for a bit?

OpenStudy (chmvijay):

why???

OpenStudy (chmvijay):

u have another question?

OpenStudy (anonymous):

i will have another one. i found this chapter a little difficult to understand.

OpenStudy (anonymous):

and the teacher SUCKS!!!!!! so we're having to self teach

OpenStudy (anonymous):

a

OpenStudy (anonymous):

right now i'm doing an open book quiz and the books isn't much help either, it really has been worse case scenario

OpenStudy (anonymous):

and our test is tomorrow...so want to make sure i can grasp these concepts

OpenStudy (anonymous):

i thank you for your help...:-)

OpenStudy (anonymous):

A sample of water releases 4.50 x 103 J of heat energy and its temperature drops from 80.0ºC to 68.0ºC. What is the mass of the water? Question 5 options: 1) 89.6g 2) 13.4g 3) 15.8g 4) 1570g

OpenStudy (anonymous):

so..solving for mass.

OpenStudy (chmvijay):

yaaaa solve it

OpenStudy (anonymous):

m=q/cpxdelta t

OpenStudy (chmvijay):

yaaa

OpenStudy (anonymous):

is this my cp 4.50 x 103 J??

OpenStudy (anonymous):

no wait..that's my heat.

OpenStudy (chmvijay):

good

OpenStudy (anonymous):

good to what ? is it cp or heat?

OpenStudy (chmvijay):

heat only as u said

OpenStudy (chmvijay):

m=q/Sxdelta t where S =4.184J for water tellme answer now

OpenStudy (anonymous):

89.627...since i have to round to sig figs..then 90g?

OpenStudy (chmvijay):

here s is called specific heat of water got it good its correct

OpenStudy (anonymous):

wait no..that's not one of the answers..so 89.6g?

OpenStudy (chmvijay):

yaaa89.6 g

OpenStudy (anonymous):

yay!! ok..

OpenStudy (anonymous):

3.17g of sodium combines with chlorine to form 8.00g of sodium chloride. What is the mass of chlorine in this sample of sodium chloride? Question 6 options: 1) 3.17g 2) 4.83g 3) 8.00g 4) 11.17g

OpenStudy (chmvijay):

enough man

OpenStudy (chmvijay):

write balanced equation

OpenStudy (anonymous):

4.83g..pellet..that as an easy one:-P

OpenStudy (anonymous):

pellet is a type-o

OpenStudy (chmvijay):

i didnt get ur question properly chlorine or chloride

OpenStudy (chmvijay):

ok may be its correct

OpenStudy (anonymous):

chlorine + sodium =sodium chloride

OpenStudy (anonymous):

i think it was asking how much chlorine was needed for sodium chloride to be 8.00

OpenStudy (anonymous):

The specific heat of aluminum is 0.900J/gºC. A 400.0g sample of aluminum at 20.0ºC absorbs 2520. J of heat energy. What is the new temperature of the aluminum? Question 9 options: 1) 7.00ºC 2) 13.0ºC 3) 20.0ºC 4) 27.0ºC

OpenStudy (chmvijay):

u solve it dude

OpenStudy (chmvijay):

assuming u dont know delta u solve i will tell u how u have to do next can u find delta T

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i'm not sure where to start..am i converting the 2520j to celsius/??

OpenStudy (chmvijay):

LOL!!!delta T=Q/M*s u do now

OpenStudy (chmvijay):

i will go dude bye do wll u know it very well

OpenStudy (anonymous):

tahnk you for your help